OpenStudy (moonlitfate):

Condense the expression to the logarithm of a single quantity. 3 ln(3) + 5 ln(z – 9)

OpenStudy (moonlitfate):

I got ln|9 (z-9)^5| Maybe I'm missing a few steps...

OpenStudy (moonlitfate):

*27 not 9

OpenStudy (anonymous):

looks good to me (with the 27 of course)

OpenStudy (moonlitfate):

It wasn't correct, and I'm not given the correct answer. Hmm..,

OpenStudy (anonymous):

get rid of the absolute value

OpenStudy (anonymous):

\[\ln(3^3)+\ln((x-9)^5)=\ln(27(x-9)^5)\] for sure maybe it was a syntax error, not enough parentheses or something

OpenStudy (moonlitfate):

I'll try to re-submit to get partial points with the parenthesis instead of absolute value things.